# Statistical Inference Leveraging Synthetic Data with Distribution-Free Guarantees

### v1

@@ Abstract
The rapid proliferation of high-quality synthetic data—generated by advanced AI models or collected as auxiliary data from related tasks—presents both opportunities and challenges for statistical inference. This paper introduces a GEneral Synthetic-Powered Inference (GESPI) framework that wraps around any statistical inference procedure to safely enhance sample efficiency by combining synthetic and real data. Our framework leverages high-quality synthetic data to boost statistical power, yet adaptively defaults to the standard inference method using only real data when synthetic data is of low quality. The error of our method remains below a user-specified bound without any distributional assumptions on the synthetic data, and decreases as the quality of the synthetic data improves. This flexibility enables seamless integration with conformal prediction, risk control, hypothesis testing, and multiple testing procedures, all without modifying the base inference method. We demonstrate the benefits of our method on challenging tasks with limited labeled data, including AlphaFold protein structure prediction, and comparing large reasoning models on complex math problems.  [^1] Software for reproducing the experiments is available at https://github.com/Meshiba/gespi.git.
@@ 1 Introduction
Statistical inference lies at the core of data-driven decision-making, enabling researchers and practitioners to draw conclusions while rigorously controlling error rates. The importance of such control cannot be overstated: uncontrolled errors can lead to misleading conclusions and costly mistakes. For example, in computational biology, researchers increasingly rely on AlphaFold’s protein structure predictions, where substantial local errors can mislead downstream applications such as drug discovery or protein design. A/B testing can be used to assess whether a new large language model (LLM) outperforms the current one—yet falsely concluding that the new model is better can lead to revenue loss or customer dissatisfaction.
A fundamental limitation of statistical inference arises when data are scarce, leading to reduced statistical power and high variability in error rates. Yet limited data is almost unavoidable in domains where data acquisition is costly, difficult, or time-consuming. For example, in protein structure prediction, experimental validation is expensive and labor-intensive, resulting in relatively few labeled structures. Similarly, comparing LLMs on complex mathematical reasoning tasks requires curating high-quality problems with verified solutions: a process that is challenging and resource-intensive.
At the same time, the availability of high-quality synthetic data presents new opportunities to mitigate data scarcity and enhance statistical efficiency. For example, such data can be generated by LLMs or diffusion models, or retrieved from related auxiliary databases. However, it is challenging to construct procedures that leverage synthetic data with provable theoretical error rate control; because synthetic data may not reflect the real-world distribution. Naively pooling synthetic and real data in standard statistical inference methods can potentially result in poor performance that depends on the (unknown) quality of the synthetic data.
This tension between opportunities and risk motivates the goal of this paper: to design a general approach that “wraps” around any statistical inference method, providing it with the ability to harness synthetic data when beneficial while attaining rigorous, distribution-free error rate control.
1.1 Preview of our method and key contributions
@@
Our framework transforms any base statistical inference method to leverage synthetic data. As in Figure˜1, let $\mathrm{Alg}$ denote the base inference method, which could represent conformal prediction, an A/B testing procedure, etc. Let $\mathcal{D}_{n}$ be the real observational dataset and $\tilde{\mathcal{D}}_{N}$ the abundant synthetic dataset, with sizes $n$ and $N\gg n$, respectively. Ideally, the synthetic data distribution would match the real one, but we make no such assumption.
@@
    Base $\mathrm{Alg}_{n,\alpha}$, uses the real data $\mathcal{D}_{n}$, targeting an error rate of $\alpha$ (e.g., 5%).
@@
    Guardrail $\mathrm{Alg}_{n,\alpha+\varepsilon}$, also uses only $\mathcal{D}_{n}$ but at a slightly higher target error rate level $\alpha+\varepsilon$ (e.g., 7%). This more relaxed level can increase the power compared to using $\alpha$.
@@
    Synthetic-powered $\mathrm{Alg}_{n,N,\alpha}$, which applies the base method to the pooled real and synthetic data, $\mathcal{D}_{n}\cup\tilde{\mathcal{D}}_{N}$, at level $\alpha$.
GESPI then aggregates (in a way we define) the outputs of these runs. We prove that this careful construction ensures the error rate never exceeds the guardrail bound $\alpha+\varepsilon$, regardless of the quality of the synthetic data (Theorems˜3.2 and 3.3). At the same time, if the synthetic dataset is well aligned with the real distribution, GESPI adapts to use the synthetic-powered outcome. In this case, our theory shows that GESPI achieves tight $\alpha$ error rate control and mimics the effect of applying the base method to a larger real dataset. Our guarantees hold in finite samples and do not require any assumptions on the synthetic data distribution.
In Section˜4 and Appendix˜C, we test the applicability of GESPI on a variety of tasks, including: (i) AlphaFold protein structure prediction, where we apply conformal risk control to bound the average fraction of residues with large prediction error; (ii) comparison of large reasoning models on math datasets, where we use hypothesis testing for win rate; (iii+iv) out-of-distribution detection, where we control the Type I error (in the single outlier case) and the family-wise error rate (for multiple outliers); (v) mechanistic interpretability of a Vision Transformer model, where we use a two-sample test to provide evidence for a property-specific role of an attention head. We also conduct ablation studies to evaluate the sensitivity of our method to $\varepsilon$ and to the synthetic data quality. These experiments support that GESPI adapts to the quality of synthetic data, providing meaningful improvements when possible while always maintaining error rate control.
@@
Figure 2: Visualization of protein structure prediction with error rate control. Panels show protein T1029 predictions with residues abstained on by (a) OnlyReal and (b) GESPI methods. Red: residues abstained on; Blue: accepted residues. Gray: real experimental structure, aligned with AlphaFold2 predicted structure. Quantitative results {abstention ratio, risk}: OnlyReal – $100\%$, $0\%$}; GESPI – $85.6\%$, $\approx 7\%$}. See text in Section˜4.1 for more details.
@@ 2 Related work
Our proposed GESPI framework is inspired by Synthetic-Powered Predictive Inference (SPI) (Bashari et al., 2025a), a recent procedure for incorporating synthetic data into conformal prediction. SPI aims to construct prediction sets with distribution-free, finite-sample coverage guarantees that hold regardless of synthetic data quality. While we share the underlying motivation of leveraging synthetic data under rigorous error rate control, our GESPI approach reformulates SPI, offering a new interpretation that serves as a foundation for addressing more general statistical inference problems. Indeed, a key distinction between the two methods is that, instead of directly modifying the mechanism for constructing prediction sets (via transportation of non-conformity scores), GESPI treats conformal prediction as a “black-box” statistical method, without altering its inner workings.
Boyeau et al. (2025); Fisch et al. (2024); Oosterhuis et al. (2024) represent another related line of research, offering methods that use a large set of synthetically generated labels, together with a small set of human labels, for unbiased model evaluation—including the construction of confidence intervals for model performance. See Chatzi et al. (2024) for ranking, and Angelopoulos et al. (2023a, b) for constructing confidence bounds for parameters of interest. Notably, this line of work assumes that the covariates of the unlabeled and labeled data are i.i.d., in striking contrast to GESPI.
We refer the reader to Appendix˜A for additional related work.
3 General synthetic-powered inference (GESPI)
@@ 3.1 GESPI for conformal prediction and risk control
Consider a predictive inference task where we are given $n$ i.i.d. (real) data points $\smash{\mathcal{D}_{n}={(X_{i},Y_{i})}_{i=1}^{n}\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}P}$, where $X_{i}\in\mathcal{X}$ and $Y_{i}\in\mathcal{Y}$ represent the features (e.g., an image) and labels (e.g., pedestrian), respectively. Given a new test input $X_{n+1}$, we aim to construct a prediction set $\smash{\widehat{C}_{n}(X_{n+1})}$ that contains the unknown test label $Y_{n+1}$, such that for a user-specified level $\alpha\in(0,1)$, e.g., 10%, the following holds:
$\mathbb{P}_{{\mathcal{D}_{n}\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}P_{X,Y},(X_{n+1},Y_{n+1})\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}P_{X,Y}}}\left\{{Y_{n+1}\notin\widehat{C}_{n}(X_{n+1})}\right\}\leq\alpha.$
Conformal prediction (Saunders et al., 1999; Vovk et al., 1999, 2005; Papadopoulos et al., 2002) is a framework that takes as input a holdout dataset $\mathcal{D}_{n}$ and transforms the output of any machine learning model into a prediction set $\widehat{C}_{n}(X_{n+1})$ for a new test point. Importantly, this prediction set satisfies the above finite-sample coverage guarantee, regardless of the underlying data distribution.
Conformal risk control extends conformal prediction beyond miscoverage rate control (0-1 loss) to any monotone risk function, such as the false negative rate or the F1 score. This framework constructs prediction sets $\widehat{C}_{n}$ with the following distribution-free risk control guarantee, for a target risk level $\alpha>0$, e.g., 10% false negative rate, and $\ell(\cdot,\cdot)$ is the loss of interest:
$\mathbb{E}_{\mathcal{D}_{n}\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}P_{X,Y},(X_{n+1},Y_{n+1})\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}P_{X,Y}}\left[\ell({\widehat{C}_{n}(X_{n+1}),Y_{n+1}})\right]\leq\alpha.$ (1)
@@
The GESPI conformal prediction set for a test point $X_{n+1}=x$ is given by $\small\hat{C}^{\textsc{GESPI}}(x):=\widehat{C}_{n,\alpha}(x)\cap\!\bigl(\widehat{C}_{n,N,\alpha}(x)\cup\widehat{C}_{n,\alpha+\varepsilon}(x)\bigr).$ (2)
The intuition behind our GESPI construction is as follows. If the synthetic and real data distributions are identical, the second term $\widehat{C}_{n,N,\alpha}(x)$ amounts to applying conformal prediction on a larger real dataset, achieving the target risk level $\alpha$ while attaining tighter and more stable prediction sets. [^2] We usually have $\widehat{C}_{n,\alpha+\varepsilon}\subseteq\widehat{C}_{n,\alpha}$, as sets get wider when a tighter error guarantee (smaller $\alpha$) is required. In this case, we have $\widehat{C}_{n,\alpha+\varepsilon}\subseteq\hat{C}^{\textsc{GESPI}}\subseteq\widehat{C}_{n,\alpha}$. When the synthetic data has high quality, we expect that $\widehat{C}_{n,N,\alpha}$ is small, and does not increase $\widehat{C}_{n,\alpha+\varepsilon}$ by much. In such a setting, we will have that $\hat{C}^{\textsc{GESPI}}$ is close to $\widehat{C}_{n,\alpha+\varepsilon}$, which can be a much tighter set than the original set $\widehat{C}_{n,\alpha}$. This explains how GESPI can lead to tighter sets. At the same time, if the synthetic data is of poor quality, there are two guardrail bounds: (i) $\widehat{C}_{n,\alpha+\varepsilon}$ ensures the risk level of $\hat{C}^{\textsc{GESPI}}$ does not exceed $\alpha+\varepsilon$; and (ii) $\widehat{C}_{n,\alpha}$ prevents the GESPI prediction set from being even wider than the base method $\widehat{C}_{n,\alpha}$.
@@ 3.2 GESPI for one-sided hypothesis testing
We now turn to describe how GESPI can be used to enhance sample efficiency in the canonical problem of one-sided hypothesis testing; see e.g., Lehmann and Romano (2005b) for an overview of hypothesis testing. Consider a parameter $\theta$ of interest. Our goal is to test the following hypothesis
$\mathcal{H}_{0}:\theta=\theta_{0}\text{ versus }\mathcal{H}_{1}:\theta>\theta_{0},$
where $\mathcal{H}_{0}$ is the null hypothesis and $\mathcal{H}_{1}$ denotes the alternative. The parameter $\theta$ could represent, for example, the prediction error of a model, or the win rate of model A compared to model B. In the latter, rejecting the null that $\theta_{0}=0.5$ provides evidence that model A outperforms model B, an application we revisit later in the experiments (Section˜4.2).
Let $\mathcal{D}_{n}=(X_{i})_{i=1}^{n}$ denote the real dataset, with $\mathcal{D}_{n}\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}P_{\theta}$, where $P_{\theta}$ is a distribution that depends on [^3] This distribution could also depend on other parameters, which are omitted here for clarity. $\theta\in\Theta$. When only a small dataset is available, statistical tests may suffer from low power; for example, the empirical mean estimate $\hat{\theta}$ can be noisy when evaluated on limited data, making it difficult to detect whether $\theta>\theta_{0}$ when using $\hat{\theta}$ as a test statistic.
@@
Intuitively, GESPI for hypothesis testing works as follows. We first apply the test $\phi_{n,N,\alpha}$ to the pooled dataset at level $\alpha$. If it rejects the null, we do not immediately reject, since the synthetic data may come from a distribution that differs significantly from the real one. To account for this, we also run the test on the real dataset $\phi_{n,\alpha+\varepsilon}$ at a slightly relaxed level $\alpha+\varepsilon$, and reject the null only if this test also rejects. In any case, if the base test $\phi_{n,\alpha}$ on the real dataset at level $\alpha$ rejects the null, we reject it immediately. This ensures that GESPI never loses power compared to the base test at level $\alpha$.
@@
Importantly, for both Type I error rate control and power, the synthetic data do not need to follow the exact distribution of the real data. This flexibility stems from the structure of one-sided tests. To see this, consider for illustration a simple setting where power is increasing [^4] This holds generally, for one-dimensional families of probability distributions with the monotone likelihood ratio property, including exponential families such as the normal mean; see Lehmann and Romano (2005b). in the true parameter $\theta$ of the real distribution. Suppose that the pooled distribution can be described by the parameter $\theta^{\text{pool}}$. To control the Type I error when the null is true, it suffices that $\theta^{\text{pool}}\leq\theta=\theta_{0}$, even if the synthetic distribution differs from the real one (i.e., $\theta^{\text{pool}}\neq\theta$). Analogously, under the alternative $\theta>\theta_{0}$, it suffices that $\theta^{\text{pool}}>\theta_{0}$, without requiring $\theta^{\text{pool}}=\theta$. This property greatly expands the range of useful synthetic data that GESPI can leverage to improve power in one-sided hypothesis testing.
3.3 GESPI for additional tasks: outlier detection and multiple testing
GESPI can be used for a number of additional statistical inference problems, including outlier detection and multiple hypothesis testing. Due to space limitations, we are only able to present a high-level overview of these here, and defer details to the appendix (See Appendices˜D and H).
@@ 3.3 GESPI for additional tasks: outlier detection and multiple testing
Outlier detection. Given a set of inliers $\mathcal{D}_{n}=(X_{i})_{i=1}^{n}$, with $X_{i}\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}P$, and a test point $X_{n+1}$, the goal is to determine whether $X_{n+1}$ is an inlier sampled from PP—or an outlier sampled from a different distribution. Formally, this can be framed as testing the null hypothesis: $\mathcal{H}_{0}:X_{n+1}\sim P$. Conformal outlier/anomaly detection Vovk et al. (2005); Bates et al. (2023); Laxhammar and Falkman (2011) provides a distribution-free test with finite-sample Type I error rate control. The associated test can then be used precisely as in Section 3.2 with GESPI to improve power with synthetic data.

### v2

@@ Abstract
The rapid proliferation of high-quality synthetic data—generated by advanced AI models or collected as auxiliary data from related tasks—presents both opportunities and challenges for statistical inference. This paper introduces a GEneral Synthetic-Powered Inference (GESPI) framework that wraps around any statistical inference procedure to safely enhance sample efficiency by combining synthetic and real data. Our framework leverages high-quality synthetic data to boost statistical power, yet adaptively defaults to the standard method using only real data when synthetic data are of low quality. The error rate of our method remains below a user-specified bound without any distributional assumptions on the synthetic data, and decreases as the quality of the synthetic data improves. This flexibility enables seamless integration with conformal prediction, risk control, hypothesis testing, and multiple testing procedures, all without modifying the base inference method. We demonstrate the benefits of our method on challenging tasks with limited labeled data, including AlphaFold protein structure prediction, and comparing large reasoning models on complex math problems. [^1] Software for reproducing the experiments is available at https://github.com/Meshiba/gespi.
@@ 1 Introduction
Statistical inference lies at the core of data-driven decision-making, enabling researchers and practitioners to draw conclusions while rigorously controlling error rates or risks. The importance of such control cannot be overstated: uncontrolled errors can lead to misleading conclusions and costly mistakes. For example, in computational biology, researchers increasingly rely on AlphaFold’s protein structure predictions, where substantial local errors can mislead downstream applications such as drug discovery or protein design. A/B testing can be used to assess whether a new large language model (LLM) outperforms the current one—yet falsely concluding that the new model is better can lead to revenue loss or customer dissatisfaction.
A fundamental limitation of statistical inference arises when data are scarce, leading to reduced utility (e.g., statistical power) and high variability in error rates. Yet limited data are almost unavoidable in domains where data acquisition is costly, difficult, or time-consuming. For example, in protein structure prediction, experimental validation is expensive and labor-intensive, resulting in relatively few labeled structures. Similarly, comparing LLMs on complex mathematical reasoning tasks requires curating high-quality problems with verified solutions: a process that is challenging and resource-intensive.
At the same time, the availability of high-quality synthetic data presents new opportunities to mitigate data scarcity and enhance statistical utility. For example, such data can be generated by LLMs or diffusion models, or retrieved from related auxiliary databases. However, it is challenging to construct procedures that leverage synthetic data with provable theoretical error rate control; because synthetic data may not reflect the real-world distribution.
To make this concrete, consider a practitioner with limited real data, abundant synthetic data, and an error-controlling algorithm, e.g., A/B testing procedure, conformal prediction, etc. How can one perform statistical inference in this setting, with the goal of obtaining informative inference while controlling the error rate?
@@
The first option is to completely ignore the synthetic data, even when they may be informative, and rely solely on the limited real data. This approach can lead to overly conservative inference, often prompting practitioners to increase the error level to increase power, at the cost of obtaining a higher error rate. The second option is to blindly trust the synthetic data as if they were real, pooling them with the limited real data. However, this approach can in general result in an uncontrolled error rate, as the synthetic data distribution can be arbitrary. Both options are insufficient.
This fundamental challenge motivates the key question of this work: how can we utilize high-quality but technically unreliable synthetic data in a reliable way? We present a general approach that “wraps” around any statistical inference method, providing it with the ability to harness synthetic data when beneficial, while attaining rigorous, distribution-free error rate control, even when the synthetic data quality is unknown a priori.
1.1 Key Contributions and Preview of Our Method
As a first step, we introduce a novel theoretical formulation for leveraging synthetic data in the context of statistical decision theory, which is based on controlling a general loss function. This formulation allows us to tackle a broad range of statistical problems in a unified framework, including hypothesis testing, multiple testing, confidence sets, etc.
Building on this formulation, our framework transforms any base statistical inference method to safely leverage synthetic data. As in Figure˜1, let $\mathrm{Alg}$ denote the base inference method, $\mathcal{D}_{n}$ the real observational dataset, and $\tilde{\mathcal{D}}_{N}$ the abundant synthetic dataset, with sizes $n$ and $N\gg n$, respectively. Ideally, the synthetic data distribution would match the real one, but we make no such assumption.
@@ 1.1 Key Contributions and Preview of Our Method
    Base $\mathrm{Alg}_{n,\alpha}$, uses the real data $\mathcal{D}_{n}$, targeting a user-specified error rate level of $\alpha$ (e.g., 5%).
@@
    Guardrail $\mathrm{Alg}_{n,\alpha+\varepsilon}$, also uses only $\mathcal{D}_{n}$ but at a slightly higher user-specified level $\alpha+\varepsilon$ (e.g., 7%).
@@
    Synthetic-powered $\mathrm{Alg}_{n,N,\alpha}$, which applies the base method to the pooled real and synthetic data $\mathcal{D}_{n}\cup\tilde{\mathcal{D}}_{N}$ at level $\alpha$.
GESPI then aggregates (in a way we define) the outputs of these runs. In terms of statistical validity, we prove in Theorems˜3.3 and F.2 that this careful construction ensures GESPI’s error rate never exceeds the guardrail bound $\alpha+\varepsilon$, even when the synthetic data are of poor quality. Relaxing to $\alpha+\varepsilon$ is essential to allow us to potentially benefit from additional synthetic data (Proposition˜E.1). At the same time, as the quality of the synthetic data improves—better matching the distribution of the gold-standard real data—the error rate can approach the desired level $\alpha$. From the perspective of statistical efficiency, we prove in Theorem˜3.4 that GESPI’s utility (e.g., power, set size) is always at least that of the base method and is bounded above by the guardrail. Our guarantees hold in finite samples and do not require any assumptions on the synthetic data distribution.
As a result, one can interpret the user-specified $\varepsilon$ as the “admission cost” for the opportunity to benefit from synthetic data in statistical inference. In the worst case (poor synthetic data), the user “pays” an additional $\varepsilon$ in error rate. In the best case (perfect synthetic data), the user enjoys enhanced sample efficiency and improved utility—where this gain increases with $\varepsilon$.
An important feature of GESPI is its ability to automatically adapt to the quality of the synthetic data. This leads to various potential benefits in sample efficiency that go beyond utility improvements, depending on the application. In this work, we highlight two main benefits of our method: (i) for tasks such as conformal prediction and risk control, GESPI reduces the variance of the empirical error rate compared to using only real data; and (ii) for hypothesis testing problems, it improves the ROC curve relative to methods that ignore the synthetic data.
To illustrate these benefits, we evaluate GESPI across a variety of tasks in Sections˜4 and C, including: (i) image classification, where we control the coverage using conformal prediction; (ii) AlphaFold protein structure prediction, where we apply conformal risk control to bound the average fraction of residues with large prediction error; (iii+iv) out-of-distribution detection, where we control the Type I error (in the single outlier case) and the family-wise error rate (for multiple outliers); (v+vi) comparison of large reasoning models on math datasets, where we use hypothesis testing for win rate, for both pairwise comparisons between models and hyperparameter selection; (vii) mechanistic interpretability of a Vision Transformer model, where we use a two-sample test to provide evidence for a property-specific role of an attention head. We also conduct ablation studies to evaluate the sensitivity of our method to $\varepsilon$ and to the synthetic data quality. These experiments support that GESPI adapts to the quality of synthetic data, providing meaningful improvements when possible while always maintaining error rate control.
@@
Figure 2: Visualization of protein structure prediction with error rate control. Panels show protein T1029 predictions with residues abstained on by (a) OnlyReal and (b) GESPI methods. Red: residues abstained on; Blue: accepted residues. Gray: real experimental structure, aligned with AlphaFold2 predicted structure. Quantitative results {abstention ratio, risk}: OnlyReal – $100\%$, $0\%$}; GESPI – $85.6\%$, $\approx 7\%$}. See text in Section˜4.2.
@@ 2 Related Work
Our proposed GESPI framework is inspired by Synthetic-Powered Predictive Inference (SPI) (Bashari et al., 2025a), a recent procedure for incorporating synthetic data into conformal prediction. SPI constructs prediction sets with distribution-free, finite-sample coverage guarantees that hold regardless of synthetic data quality. Its core idea is to modify the construction of the prediction set by applying a novel transportation to the non-conformity scores from the real space to the synthetic one.
While we share the underlying motivation of leveraging synthetic data under rigorous error rate control, our GESPI approach offers a new perspective on how to leverage synthetic data, that serves as a foundation for addressing more general statistical inference problems. Indeed, a key distinction between the two methods is that, instead of directly modifying the mechanism for constructing prediction sets—whose extension to broader statistical problems is unclear—GESPI treats the statistical inference method as a “black-box” procedure, without altering its inner workings. Accordingly, while SPI applies to conformal prediction, it does not extend to the broader statistical inference tasks considered in this paper. Additional review of works broadly related to our work is provided in Appendix˜A. Notably, these methods do not provide distribution-free control of risk when leveraging arbitrary synthetic data, which is the goal of this work.
3 General Synthetic-Powered Inference
@@ 3.1 GESPI for Conformal Prediction and Risk Control
Consider a predictive inference task where we are given $n$ i.i.d. (real) data points $\smash{\mathcal{D}_{n}={(X_{i},Y_{i})}_{i=1}^{n}\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}P}$, where $X_{i}\in\mathcal{X}$ and $Y_{i}\in\mathcal{Y}$ represent the features (e.g., an image) and labels (e.g., pedestrian), respectively. Given a new test input $X_{n+1}$, we aim to construct a prediction set $\smash{\widehat{C}_{n}(X_{n+1})}$ such that the expected loss (i.e., risk) of $\widehat{C}_{n}$ does not exceed a user-specified level $\alpha$, for a loss $\ell$ of interest:
$\mathbb{E}_{\mathcal{D}_{n}\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}P_{X,Y},(X_{n+1},Y_{n+1})\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}P_{X,Y}}\left[\ell({\widehat{C}_{n}(X_{n+1}),Y_{n+1}})\right]\leq\alpha.$
Conformal prediction (Saunders et al., 1999; Vovk et al., 1999, 2005; Papadopoulos et al., 2002) is a framework that takes a holdout dataset $\mathcal{D}_{n}$ and transforms the output of any machine learning model into a prediction set $\widehat{C}_{n}(X_{n+1})$ that satisfies the finite-sample guarantee above for the 0-1 miscoverage loss, $\mathbb{I}\{Y_{n+1}\notin\widehat{C}_{n}(X_{n+1})\}$. Conformal risk control extends this framework to any monotone loss function, such as the false negative rate or the F1 score.
@@
The GESPI conformal prediction set when the test datapoint $X_{n+1}$ takes value $x$ is given by $\small\hat{C}^{\textsc{GESPI}}(x):=\widehat{C}_{n,\alpha}(x)\cap\!\bigl(\widehat{C}_{n,N,\alpha}(x)\cup\widehat{C}_{n,\alpha+\varepsilon}(x)\bigr).$ (1)
The intuition behind our GESPI construction is as follows. If the synthetic and real data distributions are identical, the second term $\widehat{C}_{n,N,\alpha}(x)$ amounts to applying conformal prediction on a larger real dataset, achieving the target risk level $\alpha$ while attaining tighter and more stable prediction sets. [^2] We usually have $\widehat{C}_{n,\alpha+\varepsilon}\subseteq\widehat{C}_{n,\alpha}$, as sets get wider when a tighter error guarantee (smaller $\alpha$) is required. In this case, we have $\widehat{C}_{n,\alpha+\varepsilon}\subseteq\hat{C}^{\textsc{GESPI}}\subseteq\widehat{C}_{n,\alpha}$. When the synthetic data has high quality, we expect that $\widehat{C}_{n,N,\alpha}$ is small, and does not increase $\widehat{C}_{n,\alpha+\varepsilon}$ by much. In such a setting, we will have that $\hat{C}^{\textsc{GESPI}}$ is close to $\widehat{C}_{n,\alpha+\varepsilon}$, which can be a much tighter set than the original set $\widehat{C}_{n,\alpha}$. This explains how GESPI can lead to tighter sets. At the same time, if the synthetic data are of poor quality, there are two guardrail bounds: (i) $\widehat{C}_{n,\alpha+\varepsilon}$ ensures the risk level of $\hat{C}^{\textsc{GESPI}}$ does not exceed $\alpha+\varepsilon$; and (ii) $\widehat{C}_{n,\alpha}$ prevents the GESPI prediction set from being even wider than the base method $\widehat{C}_{n,\alpha}$.
@@ 3.2 GESPI for One-Sided Hypothesis Testing
We now turn to describe how GESPI can enhance sample efficiency in the canonical problem of one-sided hypothesis testing; see e.g., Lehmann and Romano (2005b) for an overview of hypothesis testing. Consider a parameter $\theta$ of interest. Our goal is to test the following hypothesis
$\mathcal{H}_{0}:\theta\leq\theta_{0}\text{ versus }\mathcal{H}_{1}:\theta>\theta_{0},$
where $\mathcal{H}_{0}$ is the null hypothesis and $\mathcal{H}_{1}$ the alternative. The parameter $\theta$ could represent, for example, the prediction error of a model, or the win rate of model A compared to model B. In the latter, rejecting the null that $\theta_{0}\leq 0.5$ provides evidence that model A outperforms model B, an application we revisit in the experiments (Section˜C.4).
Let $\mathcal{D}_{n}=(X_{i})_{i=1}^{n}$ denote the real dataset, with $\mathcal{D}_{n}\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}P_{\theta}$, where $P_{\theta}$ depends on [^3] This distribution could also depend on other parameters, which are omitted here for clarity. $\theta\in\Theta$. When only a small dataset is available, statistical tests may suffer from low power; for example, the empirical mean estimate $\hat{\theta}$ can be noisy when evaluated on limited data, making it difficult to detect whether $\theta>\theta_{0}$ when using $\hat{\theta}$ as a test statistic.
@@
Intuitively, GESPI for hypothesis testing works as follows. We first apply the test $\phi_{n,N,\alpha}$ to the pooled dataset at level $\alpha$. If it rejects the null, we do not immediately reject, since the synthetic data may come from a distribution that differs significantly from the real one. To account for this, we also run the test on the real dataset $\phi_{n,\alpha+\varepsilon}$ at a slightly relaxed level $\alpha+\varepsilon$, and reject the null only if this test also rejects. In any case, if the base test $\phi_{n,\alpha}$ on the real dataset rejects the null, we reject it immediately. This ensures that GESPI never loses power compared to the base test at level $\alpha$.
@@
Importantly, for both Type I error rate control and power, the synthetic data do not need to follow the exact distribution of the real data. This flexibility stems from the structure of one-sided tests. To see this, consider for illustration a simple setting where power is increasing [^4] This holds generally, for one-dimensional families of probability distributions with the monotone likelihood ratio property, including exponential families such as the normal mean; see Lehmann and Romano (2005b). in the true parameter $\theta$ of the real distribution. Suppose that the pooled distribution can be described by the parameter $\theta^{\text{pool}}$. To control the Type I error when the null is true, it suffices that $\theta^{\text{pool}}\leq\theta\leq\theta_{0}$, even if the synthetic distribution differs from the real one (i.e., $\theta^{\text{pool}}\neq\theta$). Analogously, under the alternative $\theta>\theta_{0}$, it suffices that $\theta^{\text{pool}}>\theta_{0}$, without requiring $\theta^{\text{pool}}=\theta$. This property greatly expands the range of useful synthetic data that GESPI can leverage to improve power in one-sided hypothesis testing.
Remark 3.1.
@@
The formulation in (2) is general and can be applied to additional hypothesis testing problems. One important problem that we revisit later in the experiments is outlier detection. Formally, given a set of inliers $\mathcal{D}_{n}=(X_{i})_{i=1}^{n}$, with $X_{i}\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}P$, and a test point $X_{n+1}$, the task is to determine whether $X_{n+1}$ is an inlier sampled from PP—or an outlier sampled from a different distribution. This can be framed as testing the null hypothesis: $\mathcal{H}_{0}:X_{n+1}\sim P$. Any test for this null—e.g. those by Vovk et al. (2005); Bates et al. (2023); Laxhammar and Falkman (2011)—can be used within our GESPI framework.
3.3 GESPI for Multiple Hypothesis Testing
GESPI can be used for a number of additional statistical inference problems. Due to space limitations, we are only able to present here a high-level overview of the multiple testing problem Lehmann and Romano (2005b) [^5] Multiple hypothesis testing has a broad range of applications across science and engineering, see for instance Benjamini and Hochberg (1995); Efron (2012); Bretz et al. (2016), etc., and defer details to the appendix (See Appendices˜D and H).